# Convective Space-Time Chaos as a Dynamical Model of Deterministic and Stochastic Turbulence

### v1

@@ Abstract
Recently, a concept of deterministic and stochastic turbulence has been introduced based on experiments with a boundary layer. A deterministic property means that identical random perturbations at the inlet lead to identical patterns downstream; during this stage, the non-identical, stochastic component grows and eventually dominates the flow further downstream. We argue that these properties can be explained by exploring the concept of convective space-time chaos, where secondary perturbations on top of a chaotic state grow but move away in the laboratory reference frame. We illustrate this with two simple models of convective space-time chaos, one is a partial differential equation describing waves on a film flowing down a plate, and the other is a set of unidirectionally coupled ordinary differential equations. To prove convective space-time chaos, we calculate the profiles of the convective Lyapunov exponent. The repeatability of the turbulent field in different identical experimental runs corresponds to the reliability of stable dynamical systems in response to random forcing. The onset of the stochastic component is quantified with the spatial Lyapunov exponent. We demonstrate how an effective randomization of the field is observed when the driving is quasiperiodic. Furthermore, we discuss space-time duality, which links sensitivity to boundary conditions in the convective space-time chaos to the usual sensitivity to initial conditions in a standard chaotic regime.
@@ I Introduction
Since the seminal papers of Lorenz [1] and Ruelle and Takens [2], chaos serves as a basis for the explanation of the deterministic unpredictability of turbulent flows. Although for flows in constrained geometries already simple low-dimensional chaotic systems provide a good description, for more developed turbulence, one needs rather advanced dynamical models of space-time chaos [3]. Although such models do not reproduce detailed features of turbulence, available only in thorough direct numerical simulations, they serve as useful conceptual tools for understanding such qualitative features as predictability, controllability, cascades across the scales, etc.
In a series of papers [4, 5, 6, 7], Borodolun, Kachanov, and co-workers have introduced, based on experiments with a boundary layer turbulence transition, the notions of deterministic and stochastic turbulence. In these experiments, random perturbations to the otherwise laminar boundary layer have been introduced in a controlled manner close to the inlet. Further downstream evolution was natural. The following features were observed:
@@
    Here, exactly the same random forcing signal was used in two different experiments. As a result, the measured turbulent fields in the two repetitions were very close to each other in a relatively large domain downstream of the forcing; this domain included both an initial region of perturbation growth and a part of the region with universal statistical properties, as described in feature 1. For example, in Fig. (3) of [7], the authors present plots of two instantaneous fields of velocity perturbations measured in two experiments, one in the morning and one in the evening, with the same random force; these two fields nearly coincide. This regime is called deterministic turbulence.
@@
    The coincidence (measured by a cross-correlation of the signals) of two fields in two experiments, as described in feature 2, is not perfect. The uncorrelated component grows exponentially downstream (see Fig. (27) of [7]), and eventually, at large distances from the inlet, correlations disappear. This regime is called stochastic turbulence.
The goal of this paper is to give an interpretation of these features in terms of deterministic models of space-time chaos. The main idea is that the described features are typical for a convective space-time chaos, where instability of the chaos is of convective nature - perturbations grow exponentially but are advected downstream, so that there is no instability at any fixed position of the system (this property of the secondary perturbations on top of chaos is analogous to the convective instability of laminar flows and wave systems [8]). Many properties of the convective space-time chaos have been already discussed in the literature [9, 10, 11, 12, 13, 14, 15, 16]. However, the reproducibility (feature 2 above) has not been mentioned; we will argue below that it corresponds to another phenomenon known in nonlinear dynamics – reliability of randomly forced systems [17].
We do not aim at a computationally extremely challenging modeling of a boundary layer transition, but instead present in Section II two tractable models of convective space-time chaos, one as a system of ordinary differential equations, and another as a nonlinear partial differential equation. In Section III we present the main tool of the analysis, the convective Lyapunov exponent. Section IV describes the same setup of a random driving at the inlet as one adopted in experiments [4, 5, 6, 7]. We describe how the features of deterministic and stochastic turbulence manifest themselves for the models of Section II. In particular, we quantify the transition from deterministic to stochastic turbulence with the spatial Lyapunov exponent. Section V is devoted to regular forcing. Especially for quasiperiodic driving (which is also feasible in experiments), one observes a nontrivial growth in signal complexity and an approach to effective “randomness”. In Section VI we discuss spatio-temporal duality in convective space-time chaos, according to which one can exchange the meaning of space and time variables, to get better insight into the features of turbulence. We conclude with a discussion in Section VII.
II Dynamical models for DST
@@ II.1 A partial derivative equation
As a basic continuous in space and time model for convective chaos, we use the Kuramoto-Sivashinsky (KS) equation. This equation was derived by Sivashinsky in 1977 for the instability of propagating flame fronts [18] and by Kuramoto in 1976-78 for the phase dynamics beyond Benjamin-Fair-type instability in active nonlinear medium [19, 20]. Previously, this equation was derived in 1974 by Nepomnyashchy [21] and Homsy [22] for the nonlinear stage of instability of films flowing down an inclined plate, and in 1975 for plasma waves [23]. The problem of waves on a flowing down film is mostly close to the boundary layer experiments [4, 5, 6, 7] (in fact, falling film can be considered as a boundary layer with an open surface) and will be adopted for the interpretation of the KS equation below. Correspondingly, we will refer to the basic equation as the Nepomnyashchy-Homsy-Kuramoto-Sivashinsky equation (NHKS).
@@
Here, $u(x,t)$ is the dimensionless deviation of the film width from the constant value, $x$ and $t$ are dimensionless space and time. Terms with higher derivatives in $x$ describe linear instability of perturbations. The lowest-order nonlinear term is responsible for the transformation of perturbations from the unstable to the stable part of the spectrum, which is sufficient for the saturation of the instability.
@@
$V=\frac{3(5^{3/2})\gamma^{1/2}G^{-1/3}\cos\alpha}{(2G\cos^{2}\alpha-\sin\alpha)^{3/2}},\quad G=\frac{ga_{0}^{2}}{\nu^{2}},\quad\gamma=\frac{\sigma}{\nu^{4/3}g^{1/3}\rho}\;.$ (2)
Here $\alpha$ is the angle of the plate to the vertical direction, $G$ and $\gamma$ are the Galilei and the Kapitza numbers; $g$ is the gravity acceleration, $\nu$ is the kinematic viscosity, $\rho$ is the density of fluid, $\sigma$ is the surface tension coefficient, $a_{0}$ is the depth of the unperturbed film. Equation (1) is written for the case of instability, the threshold of which is defined by the condition $G>G_{*}=\frac{5\sin\alpha}{2\cos^{2}\alpha}$. The NHKS equation is valid close to the instability threshold; thus, according to (2), velocity $V$ is large. However, for the sake of easy computations, below we consider not so large values of $V$; qualitatively all the effects remain the same.
@@
$\begin{gathered}\frac{du_{k}}{dt}+(6h)^{-1}[(u_{k-1}+u_{k}+u_{k+1}+3V)(u_{k+1}-u_{k-1})]+\\ h^{-2}(u_{k-1}-2u_{k}+u_{k+1})+h^{-4}(u_{k-2}-4u_{k-1}+6u_{k}-4u_{k+1}+u_{k+2})=0\;.\end{gathered}$ (3)
Below, we solve this ODE system using the Runge-Kutta 4th-order method with time step $2\cdot 10^{-4}$ and $h=0.2$.
Equation (1) should be supplemented with boundary and initial conditions. It is natural to set the inlet of the film to $x=0$ and consider the downstream domain $0\leq x<\infty$. Then the equation (1) should be supplemented by the initial condition $u(0\leq x<\infty,0)$ and the boundary conditions at $x=0$. If one considers a finite domain $0\leq x\leq L$, then also boundary conditions at $x=L$ are needed. Physically, these conditions should describe external forcing at the inlet $x=0$ and no-reflection at $x=L$. Because the NHKS equation contains higher spatial derivatives, the boundary conditions should include some conditions on derivatives. For the discrete version (3), to solve the equations for $1\leq k\leq N$, we have to define $u_{0},u_{-1},u_{N+1},u_{N+2}$. At the inlet we adopted $u_{0}=f(t)$, $u_{-1}=u_{1}$, where $f(t)$ is the external forcing of the system. At the outlet we adopted $u_{N+1}=2u_{N}-u_{N-1}$ and $u_{N+2}=2u_{N}-u_{N-2}$. This corresponds to the vanishing of the second and fourth discrete derivatives. Although this does not ensure a perfect non-reflection, distortions are restricted to a relatively narrow boundary domain.
@@ II.2 A chain of nonlinear amplifiers
The PDE model of Section II.1 is continuous in space and time. One can reduce complexity and numerical efforts by considering discrete space and continuous time, by formulating a set of ordinary differential equations (ODEs) that demonstrate features similar to those of PDE (1). We consider a “chain of nonlinear amplifiers” [11, 12]
$\frac{du_{n}}{dt}$ $=-u_{n}+F(u_{n-1}),\quad n=2,3,\ldots,N\;,\quad F(u)=1-2u^{2}\;,$ (5)
$\frac{du_{1}}{dt}$ $=-u_{n}+f(t)\;.$
Here, each unit is a low-pass linear filter, driven by the transformed field of the left neighbor. As a transformation function, we use the logistic map $F$. The first unit is driven by a given external force $f(t)$. The asymmetric in space coupling describes the propagation of perturbations from small to larger values of $n$. Equations (5) should be supplemented with initial conditions $u_{n}(0)$. If these values are close to zero, and the force is bounded $|f(t)|<1$, then the solutions at all $n$ remain in this interval $|u_{n}|\leq 1$.
Similar to the PDE of Section II.1, the chain (5) can be considered without forcing but with periodic boundary conditions $u_{0}(t)=u_{N}(t)$. Then, the only parameter is the system size $N$. For $N\geq 6$, typically, chaos is observed.
@@ II.3 A coupled map lattice
$u(j+1,n)=(1-\epsilon)F(u(j,n))+\epsilon F(u(j,n-1)),\quad n=1,2,\ldots;\quad u(j,0)=f(j)\;.$ (6)
@@ III Temporal and convective Lyapunov exponents
Lyapunov exponents are the main theoretical and numerical tool for characterization of chaos via quantification of sensitive dependence of the dynamics on initial conditions [26]. In spatio-temporal dynamics, in the case of an infinite in space medium, the LEs are not well-defined, because different norms used to characterize the size of perturbations are not equivalent. One way to circumvent this is to consider large finite systems with periodic boundary conditions. Then, the usual largest LE can be determined, which converges to a definite value as the length of the system goes to infinity [27, 26]. To distinguish this usual LE from other types, we will call it temporal LE.
@@
Practically, finite time intervals $T$ and thus also finite domains are considered, so that there is no issue of norms’ non-equivalence. Two values of the convective LE are of special importance. The maximal value $\lambda_{max}$ is the same as the temporal LE in a large spatial domain with periodic boundary conditions. The value at zero velocity $\lambda(0)$ is the temporal LE in a large spatial domain with non-reflecting (open) boundary conditions.
@@
Numerics of the convective exponent (Fig. 2) yields the best fit with slightly distorted expression (13), as indicated in the caption. The usual LE here is $\lambda(0)=-1$. This result is evident, because the linear matrix in (10) is a lower triangular matrix with all diagonal elements equal to 1-1. The largest LE $\lambda_{max}\approx 0.41$ is achieved at $v\approx 1.55$ and is positive. Thus, the system belongs to the class of convective space-time chaos.

### v2

@@ Abstract
Recently, a concept of deterministic and stochastic turbulence has been introduced based on experiments with a boundary layer. In these experiments, the flow was driven with controlled random perturbation; in addition, natural ambient noise was also present. Deterministic property manifested itself as repeatability of turbulence patterns induced by identical random perturbations at the inlet (deterministic turbulence). A stochastic non-identical component originating from natural noise grows and eventually dominates the flow further downstream (stochastic turbulence). We argue that these properties can be explained by exploring the concept of convective space-time chaos, where secondary perturbations on top of a chaotic state grow but move away in the laboratory reference frame. We illustrate this with two simple models of convective space-time chaos, one is a partial differential equation describing waves on a film flowing down a plate, and the other is a set of unidirectionally coupled ordinary differential equations. To prove convective space-time chaos, we calculate the profiles of the convective Lyapunov exponent. The repeatability of the turbulent field in different identical experimental runs corresponds to the reliability of stable dynamical systems in response to random forcing. The onset of the stochastic component is quantified with the spatial Lyapunov exponent. We demonstrate how an effective randomization of the field is observed when the driving is quasiperiodic. Furthermore, we discuss space-time duality, which links sensitivity to boundary conditions in the convective space-time chaos to the usual sensitivity to initial conditions in a standard chaotic regime.
@@ I Introduction
Since the seminal papers of Lorenz [1] and Ruelle and Takens [2], chaos serves as a basis for the explanation of the deterministic unpredictability of turbulent flows. Although for flows in constrained geometries already simple low-dimensional chaotic systems provide a good description, for more developed turbulence, one needs rather advanced dynamical models of space-time chaos [3]. Albeit such models do not reproduce the detailed features of turbulence, which are only available through thorough direct numerical simulations, they serve as useful conceptual tools for understanding qualitative features such as predictability, controllability, cascades across scales, etc.
An important aspect of the space-time chaos and turbulence is the nature of the linear instability of a steady state (of a laminar flow). According to the classical stability theory [4, 5, 6] one distinguishes two types of instability in homogeneous nonequilibrium media: absolute and convective. In convectively unstable systems, a local in space perturbation grows but moves away from the place of its initiation, so that the perturbation in the region where it was imposed decays. In absolutely unstable systems, a perturbation grows at the place where it was imposed. Convective instability appears in beam-plasma systems [4, 7], in electronics (e.g., in a traveling wave tube [8, 9]), in chemical reactions [10]. In hydrodynamics, convective instability appears in many open-flow systems, like wakes, capillary jets, and two-dimensional Poiseuille flow [11]. Two examples we discuss in more detail below, are sketched in Fig. 1: boundary layer flow and waves on a liquid film over an inclined plate.
Figure 1: Panel (a): a sketch of the experiments on controlled transition in a boundary layer (adapted after Figs. 1,2 in [12]). Close to the inlet region, a ribbon band is embedded into the plate, allowing for the insertion of controlled perturbations in the flow. Panel (b): a sketch of the experiments on controlled transition for waves on a thin film flow over an inclined plate (adopted from Fig. 1 in [13]). A perturbation to the flow at the inlet is performed by applying pressure variations to the entrance manifold.
A common feature of convectively unstable systems is that one cannot directly apply notions of deterministic chaos, such as sensitive dependence on initial conditions, to them. Indeed, initial perturbations are “blown away” and are not relevant at large times when observing a fixed finite domain. Instead, one needs constantly acting perturbations close to the inlet zone to have a nontrivial dynamical regime [14, 15, 16, 17, 18, 19, 20, 21, 22]. In experiments, two types of such perturbations have been explored. In many cases, chaos and turbulence develop from the natural, ambient noise. In other setups, one imposes an actuator that produces controlled perturbations. If these controlled perturbations are significantly larger than the level of natural noise, they dominate the close to the inlet domain of spatial growth and development. Typically, periodic in time perturbations are imposed to produce waves, properties of which are known from the theory, see [23, 24, 12] for such experiments with a boundary layer (Fig. 1(a)), and [25, 13, 26] for experiments with waves on liquid film (Fig. 1(b)). The resulting periodic structures can also be unstable, and natural noise leads to the loss of periodicity and the establishment of a turbulent state.
In a series of experiments [27, 28, 29, 30], Borodulin, Kachanov, and co-workers used the same experimental setup for actuation of perturbation in a boundary layer Fig. 1(a) to impose stochastic macroscopic perturbations with controllable statistical properties. In these experiments, they discovered interesting features that enabled them to introduce the concepts of “deterministic” and “stochastic” turbulence. These features appear to be common to systems with convective space-time chaos, and can be formulated independently of particular hydrodynamical aspects of the boundary layer transition. The following properties were emphasized by Borodulin, Kachanov, and co-workers:
@@
    A setup, where exactly the same random forcing signal was used in two different experiments. As a result, the measured turbulent fields in the two repetitions were very close to each other in a relatively large domain downstream of the forcing; this domain included both an initial region of perturbation growth and a part of the region with universal statistical properties, as described in feature 1. For example, in Fig. (3) of [30], the authors present plots of two instantaneous fields of velocity perturbations measured in two experiments, one in the morning and one in the evening, with the same random force; these two fields nearly coincide. This regime is called deterministic turbulence.
@@
    The coincidence (measured by a cross-correlation of the signals) of two fields in two experiments, as described in feature 2, is not perfect. The uncorrelated component (which is unavoidable in experiments due to ambient noise) grows exponentially downstream (see Fig. (27) of [30]), and eventually, at large distances from the inlet, correlations between fields in two runs disappear. This regime is called stochastic turbulence.
The goal of this paper is to give an interpretation of these features in terms of deterministic models of convective space-time chaos. In particular, we will argue that the reproducibility (feature two above) directly corresponds to another phenomenon known in nonlinear dynamics – reliability of randomly forced systems [31]. We emphasize here that our aim is not to model a boundary layer transition that is extremely challenging computationally, but rather to provide theoretical dynamical tools that can characterize the features of deterministic and stochastic turbulence as formulated above. These tools are not specific to hydrodynamics, but can be generally implemented in other manifestations of convective space-time chaos, such as in plasma physics and electronics.
We start with formulating in Section II two tractable models of convective space-time chaos, one as a system of ordinary differential equations, and another as a nonlinear partial differential equation. This partial differential equation describes waves on an inclined liquid film (Fig. 1(b)), which is a paradigmatic case of convective instability. In Section III we present the main tool of the analysis, the convective Lyapunov exponent. Section IV describes the same setup of a macroscopic random driving at the inlet as one adopted in experiments [27, 28, 29, 30]. We describe how the features of deterministic and stochastic turbulence manifest themselves for the models of Section II. In particular, we quantify the transition from deterministic to stochastic turbulence with the spatial Lyapunov exponent. Section V is devoted to regular forcing. Especially for quasiperiodic driving (which is also feasible in experiments), one observes a nontrivial growth in signal complexity and an approach to effective “randomness”. In Section VI we discuss spatio-temporal duality in convective space-time chaos, according to which one can exchange the meaning of space and time variables, to get better insight into the features of convective turbulence. We conclude with a discussion in Section VII.
II Dynamical models for convective space-time chaos
@@ II.1 A partial derivative equation
As a basic model for convective chaos continuous in space and time, we use the Kuramoto-Sivashinsky (KS) equation. This equation was derived by Sivashinsky in 1977 for the instability of propagating flame fronts [32] and by Kuramoto in 1976-78 for the phase dynamics beyond Benjamin-Fair-type instability in active nonlinear medium [33, 34]. Previously, this equation was derived in 1974 by Nepomnyashchy [35] and Homsy [36] for the nonlinear stage of instability of films flowing down an inclined plate, and in 1975 for plasma waves [37]. The problem of waves on a flowing down film is mostly close to the boundary layer experiments [27, 28, 29, 30] (in fact, falling film can be considered as a boundary layer with an open surface) and will be adopted for the interpretation of the KS equation below. Correspondingly, we will refer to the basic equation as the Nepomnyashchy-Homsy-Kuramoto-Sivashinsky equation (NHKS).
@@
Here, $u(x,t)$ represents the dimensionless deviation of the film width from its constant value; $x$ and $t$ are dimensionless space and time coordinates, respectively. Terms with higher derivatives in $x$ describe linear instability of perturbations. The lowest-order nonlinear term is responsible for the transformation of perturbations from the unstable to the stable part of the spectrum, which is sufficient for the saturation of the instability.
@@
$\begin{gathered}V=\frac{3(5^{3/2})\gamma^{1/2}G^{-1/3}\cos\alpha}{(2G\cos^{2}\alpha-\sin\alpha)^{3/2}},\quad G=\frac{ga_{0}^{2}}{\nu^{2}},\\ \gamma=\frac{\sigma}{\nu^{4/3}g^{1/3}\rho}\;.\end{gathered}$ (2)
Here $\alpha$ is the angle of the plate to the vertical direction, $G$ and $\gamma$ are the Galilei and the Kapitza numbers; $g$ is the gravity acceleration, $\nu$ is the kinematic viscosity, $\rho$ is the density of fluid, $\sigma$ is the surface tension coefficient, $a_{0}$ is the depth of the unperturbed film. Equation (1) is written for the case of instability, the threshold of which is defined by the condition $G>G_{*}=\frac{5\sin\alpha}{2\cos^{2}\alpha}$. The NHKS equation is valid close to the instability threshold; thus, according to (2), velocity $V$ is large. However, for the sake of easy computations, below we consider not so large values of $V$; qualitatively, all the effects remain the same.
@@
$\begin{gathered}\frac{du_{k}}{dt}+(6h)^{-1}(u_{k-1}+u_{k}+u_{k+1}+3V)(u_{k+1}-u_{k-1})+\\ h^{-2}(u_{k-1}-2u_{k}+u_{k+1})+\\ h^{-4}(u_{k-2}-4u_{k-1}+6u_{k}-4u_{k+1}+u_{k+2})=0\;.\end{gathered}$ (3)
Below, we solve this ODE system using the Runge-Kutta 4th-order method with time step $2\cdot 10^{-4}$ and $h=0.2$. This spatial step is small enough to ensure spatial smoothness of the solution.
Equation (1) should be supplemented with boundary and initial conditions. It is natural to set the inlet of the film to $x=0$ and consider the downstream domain $0\leq x<\infty$. Then, equation (1) should be supplemented by the initial condition $u(0\leq x<\infty,0)$ and the boundary conditions at $x=0$. If one considers a finite domain $0\leq x\leq L$, then also boundary conditions at $x=L$ are needed. Physically, these conditions should describe external forcing at the inlet $x=0$ and no-reflection at $x=L$. Because the NHKS equation contains higher spatial derivatives, the boundary conditions should include some conditions on derivatives. For the discrete version (3), to solve the equations for $1\leq k\leq N$, we have to define $u_{0},u_{-1},u_{N+1},u_{N+2}$. At the inlet we adopted $u_{0}=f(t)$, $u_{-1}=u_{1}$, where $f(t)$ is the external forcing of the system. At the outlet we adopted $u_{N+1}=2u_{N}-u_{N-1}$ and $u_{N+2}=2u_{N}-u_{N-2}$. This corresponds to the vanishing of the second and fourth discrete derivatives. Although this does not ensure a perfect non-reflection, distortions are restricted to a relatively narrow boundary domain.
@@
We conclude this section by mentioning that Eq. (1) is valid also for phase turbulence in the complex Ginzburg-Landau equation. The latter equation describes the evolution of a complex amplitude $A(x,t)$ of an unstable wave close to the instability threshold, it reads (cf. eq. (10.63) in [6])
$(\partial_{t}+s\partial_{x})A=A+(1+c_{1})\partial_{x}^{2}A-(1-ic_{3})|A|^{2}A\;.$ (5)
A spatially homogeneous solution with amplitude $|A|=1$ is stable if $1-c_{1}c_{3}>0$ (Newell’s criterion) but becomes unstable if $1-c_{1}c_{3}<0$. Close to this threshold, the real amplitude $|A|$ remains close to one, but the phase $\theta(x,t)=\text{arg}(A)$ is modulated. To describe these modulations, one reduces to a single equation for the dynamical variable $\theta$, assuming that the amplitude is slaved by the variations of the phase [34]. Such a reduction is, e.g., performed in Refs. [40, 41]; it yields
$\partial_{t}\theta+s\partial_{x}\theta-c_{3}-(1-c_{1}c_{3})\partial_{x}^{2}\theta+\frac{c_{1}^{2}}{2}\partial_{x}^{4}\theta+(c_{1}+c_{3})(\partial_{x}\theta)^{2}=0\;.$ (6)
By setting $u(x,t)=\partial_{x}\theta(x,t)$, after a proper rescaling of the variables, this equation reduces to Eq. (1).
@@ II.2 A chain of nonlinear amplifiers
The PDE model of Section II.1 is continuous in space and time. One can reduce complexity and numerical efforts by considering a discrete space and continuous time, and by formulating a set of ordinary differential equations (ODEs) that exhibit features similar to those of PDE (1). We consider a “chain of nonlinear amplifiers” [16, 17]
$\frac{du_{n}}{dt}$ $=-u_{n}+F(u_{n-1}),\quad n=2,3,\ldots,N\;,$ (7)
$\frac{du_{1}}{dt}$ $=-u_{1}+f(t)\;.$
Here, each unit is a low-pass linear filter, driven by the transformed field of the left neighbor. As a transformation function, we use the logistic map $F(u)=1-2u^{2}$. The first unit is driven by a given external force $f(t)$. The asymmetric in space coupling describes the propagation of perturbations from small to larger values of $n$. Equations (7) should be supplemented with initial conditions $u_{n}(0)$. If these values are close to zero, and the force is bounded $|f(t)|<1$, then the solutions at all $n$ remain in this interval $|u_{n}|\leq 1$.
Similar to the PDE of Section II.1, the chain (7) can be considered without forcing but with periodic boundary conditions $u_{0}(t)=u_{N}(t)$. Then, the only parameter is the system size $N$. For $N\geq 6$, typically, space-time chaos is observed.
@@ II.3 A coupled map lattice
$\begin{gathered}u(j+1,n)=(1-\epsilon)F(u(j,n))+\epsilon F(u(j,n-1)),\\ n=1,2,\ldots;\quad u(j,0)=f(j)\;.\end{gathered}$ (8)
@@ III Temporal and convective Lyapunov exponents
Lyapunov exponents (LEs) are the main theoretical and numerical tool for characterization of chaos via quantification of sensitive dependence of the dynamics on initial conditions [42]. In spatio-temporal dynamics, in the case of a spatially infinite medium, the LEs are not well-defined, because different norms used to characterize the size of perturbations are not equivalent. One way to circumvent this is to consider large finite systems with periodic boundary conditions. Then, the usual largest LE can be determined, which converges to a definite value as the length of the system goes to infinity [43, 42]. To distinguish this usual LE from other types, we will call it temporal LE.
@@
Practically, finite time intervals $T$ and thus also finite domains are considered, so that there is no issue of norms’ non-equivalence. Two values of the convective LE are of special importance. The maximal value $\lambda_{max}$ is the same as the temporal LE in a large spatial domain with periodic boundary conditions. The value at zero velocity $\lambda(0)$ is the temporal LE in a finite spatial domain with non-reflecting (open) boundary conditions.
@@
Numerics of the convective exponent (Fig. 3) yields the best fit with slightly distorted expression (15), as indicated in the caption. The usual LE here is $\lambda(0)=-1$. This result is evident because the linear matrix in (12) is a lower triangular matrix with all diagonal elements equal to 1-1. The largest LE $\lambda_{max}\approx 0.41$ is achieved at $v\approx 1.55$ and is positive. Thus, the system belongs to the class of convective space-time chaos.